OpenStudy (anonymous):

Question 5.5. Use the substitution method to solve the following system of linear equations: x + y = 6 x = –y + 2 (Points : 4) The solution is (2, 4). The solution is (4, 2). There is no solution. There are an infinite number of solutions.

OpenStudy (texaschic101):

x = -y + 2 so sub in -y + 2 in for x in the 1st equation x + y = 6 (-y + 2) + y = 6 -- combine like terms 2 = 6 (incorrect) -- no solution